OpenStudy (goformit100):

Find the smallest integer K which when divided by 6, 5, 4, 3, 2 successively leaves remainders 5, 4, 3, 2 and 1 respectively.

OpenStudy (goformit100):

@ajprincess

OpenStudy (amistre64):

chinese remainders can take awhile :)

OpenStudy (ajprincess):

R u familiar with chinese theorem? @goformit100

OpenStudy (goformit100):

No Madam I am not familiar with it.

OpenStudy (amistre64):

i recall it as\[\sum aMy\]

OpenStudy (ajprincess):

http://mathworld.wolfram.com/ChineseRemainderTheorem.html. Pls take a look at this and see if u can understand it:)

OpenStudy (amistre64):

does the CRT find the smallest? or just "a" number?

OpenStudy (goformit100):

ok let me check it out.

OpenStudy (ajprincess):

it depends on the values of \[\bar n_i\] we find. @amistre64

OpenStudy (goformit100):

I did the question as LCM of 6, 5, 4, 3, 2, 1 = 60 So Number is 60 – 1 = 59 is it correct ?

OpenStudy (goformit100):

I got 59 as the required K.

OpenStudy (amistre64):

the CRT a=remainder, M=product of divisors, divided by the required divisor; and y is a little complicated, My = 1 (mod divisor) 5 6!/6 ; y 6!/6 = 1 (mod 6) + 4 6!/5 ; y 6!/5 = 1 (mod 5) + 3 6!/4 ; y 6!/4 = 1 (mod 4) + 2 6!/3 ; y 6!/3 = 1 (mod 3) + 1 6!/2 ; y 6!/2 = 1 (mod 2)

OpenStudy (amistre64):

if 59 fits the bill, then it would be simple enough to chk

OpenStudy (goformit100):

Thank you Sir and Madam.

OpenStudy (amistre64):

59 = 60 +-1 59 = 60 +-2 these remainders dont seem to be +1, +2, +3, ...

OpenStudy (amistre64):

but i might have done that a little off

OpenStudy (amistre64):

59/6 = .9 + 5 59/5 = .11 + 4 59/4 = .14 + 3 it looks promising

OpenStudy (goformit100):

Thankyou Sir